# Quantum Inspired Excited States Calculations for Molecules Based on Contextual Subspace and Symmetry Optimizations

### v1

@@ Abstract
Quantum-inspired methods for excited-state calculations remain underexplored in Noisy Intermediate-Scale Quantum (NISQ) hardware, despite their critical role in photochemistry and material science. Here, we propose a resource-efficient framework that integrates the contextual subspace (CS) method with the Variational Quantum Deflation (VQD) algorithm to enable systematic excited-state calculations for molecules while reducing qubit requirements. On the basis of the numerical results, we find that it is unproblematic to utilize this combination in calculating the excited state to reduce qubits. Furthermore, we demonstrate that the implementation of a spin-conserving hardware-efficient ansatz, namely the $\mathcal{N}(\theta_{x},\theta_{y},\theta_{z})$ block ansatz, allows exploitation of spin symmetry within the projected subspace, thereby achieving further reductions in computational resource demands. Compared to the commonly used $R_{y}R_{z}$ ansatz, using the $\mathcal{N}(\theta_{x},\theta_{y},\theta_{z})$ ansatz can reduce the number of optimization iterations by up to 3 times at a similar circuit depth.
1 Introduction
@@
2 Contextual subspace for excited state calculation
We first give a brief introduction of Contextual Subspace Variational Quantum Eigensolver (CS-VQE) [12, 18, 19] based on some original papers. For a more detailed and rigorous explanation, see the systematic establishment by Weaving et al. [20]. After giving the foundation knowledge, we show the performance of combining the contextual subspace method with the variational quantum deflation method to calculate the excited state.
From a stabilizer perspective, the contextual subspace projection has some kind of similarity to the qubit-tapering technique [9]. It is shown that qubit tapering can be integrated by a contextual subspace projection. Thus we also present our results with the qubit tapering routine in Sec. 3. The concept of quantum contextuality originates from the Bell-Kochen-Specker [21, 22] theorem. Kirby and Love first built the condition test for an arbitrary Hamiltonian to determine noncontextuality. Contextuality subspace VQE aims at solving the ground state of a qubit Hamiltonian. This qubit Hamiltonian is composed of terms called Pauli string,
@@ 2 Contextual subspace for excited state calculation
$U_{r}gU_{r}^{{}^{\dagger}}=\sigma_{p}^{(i)},\,\forall i\in G.$ (5)
@@
3 Numerical results
3.1 CS-VQD using problem-inspired ansatz
@@
$|\psi(\bm{\theta})\rangle=U(\bm{\theta})|\psi_{ref}\rangle=e^{\hat{T}_{\text{UCCSD}}-\hat{T}^{\dagger}_{\text{UCCSD}}}|\psi_{ref}\rangle.$ (10)
@@
3.2 Application of spin conservation ansatz
@@ 4 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grant 62304037, the Natural Science Foundation of Jiangsu Province under Grant BK20230828, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant BK20233002, the Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001, the Southeast University Interdisciplinary Research Program for Young Scholars under Grant 2024FGC1005, the Fundamental Research Funds for the Central Universities under Grant 2242024K40016, and the Start-up Research Fund of Southeast University under Grant RF1028623173.

### v3

# Efficient Excited-State Calculations for Molecules Based on Contextual Subspace Method and Symmetry Optimizations

@@ Abstract
Quantum computing methods for excited-state calculations remain underexplored in Noisy Intermediate-Scale Quantum (NISQ) hardware, despite their critical role in photochemistry and material science. Herein, we propose a resource-efficient framework that integrates the contextual subspace (CS) method with the Variational Quantum Deflation (VQD) algorithm to enable systematic excited-state calculations for molecules while reducing qubit requirements. On the basis of the numerical results, we find that it is unproblematic to utilize this combination in calculating the excited state to reduce qubits. Furthermore, we demonstrate that the implementation of a spin-conserving hardware-efficient ansatz, namely the $\mathcal{N}(\theta_{x},\theta_{y},\theta_{z})$ block ansatz, allows exploitation of spin symmetry within the projected subspace, thereby achieving further reductions in computational resource demands. Compared to the commonly used $R_{y}R_{z}$ ansatz, using the $\mathcal{N}(\theta_{x},\theta_{y},\theta_{z})$ ansatz can reduce the number of optimization iterations by up to 3 times at a similar circuit depth.
preprint: APS/123-QED
I Introduction
@@
II Contextual subspace for excited state calculation
We first give a brief introduction to the Contextual Subspace Variational Quantum Eigensolver (CS-VQE) [12, 18, 19] based on some original papers. For a more detailed and rigorous explanation, see the systematic establishment by Weaving et al. [20]. After giving the foundation knowledge, we show the performance of combining the contextual subspace method with the variational quantum deflation method to calculate the excited state.
From a stabilizer perspective, the contextual subspace projection has some kind of similarity to the qubit-tapering technique [9]. It is shown that qubit tapering can be integrated by a contextual subspace projection. Thus we also present our results with the qubit tapering routine in Sec. III. The concept of quantum contextuality originates from the Bell-Kochen-Specker [21, 22] theorem. Kirby and Love first built the condition test for an arbitrary Hamiltonian to determine noncontextuality. Contextuality subspace VQE aims at solving the ground state of a qubit Hamiltonian. This qubit Hamiltonian is composed of Pauli strings which can be defined as,
@@ II Contextual subspace for excited state calculation
$U_{r}gU_{r}^{{}^{\dagger}}=\sigma_{p}^{(i)},\,\forall g\in G.$ (5)
@@
III Numerical results
III.1 CS-VQD using problem-inspired ansatz
@@
$|\psi(\vec{\theta})\rangle=U(\vec{\theta})|\psi_{ref}\rangle=e^{\hat{T}_{\text{UCCSD}}-\hat{T}^{\dagger}_{\text{UCCSD}}}|\psi_{ref}\rangle.$ (10)
@@
III.2 Application of spin conservation ansatz
@@ IV Conclusion
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grant 62304037, in part by the Natural Science Foundation of Jiangsu Province under Grant BK20230828, in part by the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant BK20233002, in part by the Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001, in part by the Southeast University Interdisciplinary Research Program for Young Scholars under Grant 2024FGC1005, in part by the Fundamental Research Funds for the Central Universities under Grant 2242024K40016, and in part by the Start-up Research Fund of Southeast University under Grant RF1028623173.